OpenStudy (anonymous):

explain how to find the solution for the equation 15+v=32

OpenStudy (anonymous):

First we need to isolate v by removing 15 from the side it is on and moving it to the other side. How would we do this?

OpenStudy (anonymous):

@Shadowreaper666 i cant understand u

OpenStudy (xxdragoonxx):

Take away 5 from both sides.

OpenStudy (anonymous):

so what i right @xxDragoonxx i know u help few days go with @runner1234568

OpenStudy (anonymous):

We need to find what we need to add to 15 to get 32. If you subtract 15 from the left side of the equation(to remove 15) and from the right side to move the 15 you will get what v equals. The first step would make the equation v = 32 - 15. Just solve from there.

OpenStudy (anonymous):

17

OpenStudy (anonymous):

Replace v in 15 + v = 32 with 17 to check your answer.

OpenStudy (anonymous):

thanks u help me alot could u help me with more problems

OpenStudy (anonymous):

what should i right in the answer then

OpenStudy (anonymous):

Sure i can help you with more problems.

OpenStudy (xxdragoonxx):

5 + -5 + v = 32 + -5 5+5-5=0 v + 0 = -5 + 32 = 27 32 + -5 = v -5 + 32 = 27 v = 27

OpenStudy (anonymous):

but still how should i write my answer

OpenStudy (anonymous):

I don't think that is correct @xxDragoonxx this should be the answer 15 + v = 32 -15 -15 v = 32 - 15 v = 17

OpenStudy (anonymous):

so that i should write in my answer box

OpenStudy (anonymous):

could u help me with more know if we are done with this problem

OpenStudy (anonymous):

You could write that in the answer box or you could add in text to explain what the next step is. The choice is yours.

OpenStudy (anonymous):

the work i willl pick could u help me with more

OpenStudy (anonymous):

problems

OpenStudy (anonymous):

Sure just post them here

OpenStudy (anonymous):

determine whether each solution is reasonable. write solution or not a solution . explain your reasoning 150-y=120;y=30 512+m=832;m=122 560 divide n=70;n=8

OpenStudy (anonymous):

these are not queation i need help

OpenStudy (anonymous):

Stop send the message

OpenStudy (anonymous):

what

OpenStudy (anonymous):

It was spamming my screen and i could not do anything

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

coudl u help me

OpenStudy (anonymous):

One min

OpenStudy (anonymous):

k srry i bothering or annoying u my bad

OpenStudy (anonymous):

150 - y = 120 ; y = 30 This one would be true because 150 - 30 = 120. Just replace the letter with the number it equals to solve for the next 2(Hint: Use a calculator). 512 + m = 832 ; m = 122 Does replacing m with 122 make the equation true? 560/n = 70 ; n = 8 Does replacing n with 8 make the equation true? Remember for these you are just trying to find if the variable makes the equation true or not.

OpenStudy (anonymous):

thats what my teacher said too

OpenStudy (anonymous):

i am not sure how to slove it

OpenStudy (anonymous):

so how would figure it out

OpenStudy (anonymous):

the secound one is true i am not sure about the last one

OpenStudy (anonymous):

are u there

OpenStudy (anonymous):

Sorry tabbed out of this and forgot to check it. Give me 2 mins

OpenStudy (anonymous):

In the second equation instead of typing m type 122 and in the third equation instead of typing n type in 8. Then use a calculator to solve the equations. I would love to help you more but i gtg good luck on your work. Post a new question if you need more help and some else can help you out.